solving by substitution method. 8m+n=3
m-9n=46

Answer by swincher4391(1107)

Let's solve for m.
Then m = 46+9n, using the 2nd equation.
Take 46+9n and place it where you see m in the 1st equation:
8(46+9n) + n = 3
Distribute.
368 + 72n +n = 3
73n = -365
n = -365/73
n+=+-5
Take n = -5 and plug it back into either ORIGINAL equation.
Let's put it back into the first.
Then:8m - 5 = 3
8m = 8
m = 8/8
m+=1
So our solution is (1,-5) or m=1n=-5